### v2

# Improving Text Embeddings with Large Language Models

Abstract
In this paper, we introduce a novel and simple method for obtaining high-quality text embeddings using only synthetic data and less than k training steps. Unlike existing methods that often depend on multi-stage intermediate pre-training with billions of weakly-supervised text pairs, followed by fine-tuning with a few labeled datasets, our method does not require building complex training pipelines or relying on manually collected datasets that are often constrained by task diversity and language coverage. We leverage proprietary LLMs to generate diverse synthetic data for hundreds of thousands of text embedding tasks across nearly languages. We then fine-tune open-source decoder-only LLMs on the synthetic data using standard contrastive loss. Experiments demonstrate that our method achieves strong performance on highly competitive text embedding benchmarks without using any labeled data. Furthermore, when fine-tuned with a mixture of synthetic and labeled data, our model sets new state-of-the-art results on the BEIR and MTEB benchmarks.
1 Introduction
Text embeddings are vector representations of natural language that encode its semantic information. They are widely used in various natural language processing (NLP) tasks, such as information retrieval (IR), question answering, semantic textual similarity, bitext mining, item recommendation, etc. In the field of IR, the first-stage retrieval often relies on text embeddings to efficiently recall a small set of candidate documents from a large-scale corpus using approximate nearest neighbor search techniques. Embedding-based retrieval is also a crucial component of retrieval-augmented generation (RAG) (Lewis et al., 2020), which is an emerging paradigm that enables large language models (LLMs) to access dynamic external knowledge without modifying the model parameters. Source attribution of generated text is another important application of text embeddings (Gao et al., 2023) that can improve the interpretability and trustworthiness of LLMs.
Previous studies have demonstrated that weighted average of pre-trained word embeddings (Pennington et al., 2014; Arora et al., 2017) is a strong baseline for measuring semantic similarity. However, these methods fail to capture the rich contextual information of natural language. With the advent of pre-trained language models (Devlin et al., 2019), Sentence-BERT (Reimers and Gurevych, 2019) and SimCSE (Gao et al., 2021) have been proposed to learn text embeddings by fine-tuning BERT on natural language inference (NLI) datasets. To further enhance the performance and robustness of text embeddings, state-of-the-art methods like E5 (Wang et al., 2022b) and BGE (Xiao et al., 2023) employ a more complex multi-stage training paradigm that first pre-trains on billions of weakly-supervised text pairs, and then fine-tunes on several labeled datasets.
Existing multi-stage approaches suffer from several drawbacks. Firstly, they entail a complex multi-stage training pipeline that demands substantial engineering efforts to curate large amounts of relevance pairs. Secondly, they rely on manually collected datasets that are often constrained by the diversity of tasks and the coverage of languages. For instance, Instructor (Su et al., 2023) is only trained on instructions from English datasets, whereas BGE (Xiao et al., 2023) only focuses on high-resource languages such as English and Chinese. Moreover, most existing methods employ BERT-style encoders as the backbone, neglecting the recent advances of training better LLMs and related techniques such as context length extension (Rozière et al., 2023).
In this paper, we propose a novel method for text embeddings that leverages LLMs to overcome the limitations of existing approaches. We use proprietary LLMs to generate synthetic data for a diverse range of text embedding tasks in languages, covering hundreds of thousands of embedding tasks. Specifically, we use a two-step prompting strategy that first prompts the LLMs to brainstorm a pool of candidate tasks, and then prompts the LLMs to generate data conditioned on a given task from the pool. To cover various application scenarios, we design multiple prompt templates for each task type and combine the generated data from different templates to boost diversity. For the text embedding models, we opt for fine-tuning powerful open-source LLMs rather than small BERT-style models. Since LLMs such as Mistral (Jiang et al., 2023) have been extensively pre-trained on web-scale data, contrastive pre-training offers little additional benefit.
We demonstrate that Mistral-7B, when fine-tuned solely on synthetic data, attains competitive performance on the BEIR (Thakur et al., 2021) and MTEB (Muennighoff et al., 2023) benchmarks. This is particularly intriguing considering that this setting does not involve any labeled data. When fine-tuned on a mixture of synthetic and labeled data, our model achieves new state-of-the-art results, surpassing previous methods by a significant margin (+). The entire training process requires less than k steps.
Moreover, we empirically validate that our model can effectively perform personalized passkey retrieval for inputs up to k tokens by altering the rotation base of the position embeddings, extending the context length beyond the conventional token limit. Regarding its multilinguality, our model excels on high-resource languages. However, for low-resource languages, there is still room for improvement as current open-source LLMs are not adequately pre-trained on them.
2 Related Work
Text Embeddings are continuous low-dimensional representations of text and have been extensively applied to various downstream tasks such as information retrieval, question answering, and retrieval-augmented generation (RAG). Early work on text embeddings includes latent semantic indexing (Deerwester et al., 1990) and weighted average of word embeddings (Mikolov et al., 2013). More recent methods exploit supervision from natural language inference (Bowman et al., 2015) and labeled query-document pairs, such as the MS-MARCO passage ranking dataset (Campos et al., 2016), to train text embeddings (Reimers and Gurevych, 2019; Conneau et al., 2017; Gao et al., 2021). However, labeled data are often limited in terms of task diversity and language coverage. To address this challenge, methods like Contriever (Izacard et al., 2021), OpenAI Embeddings (Neelakantan et al., 2022), E5 (Wang et al., 2022b), and BGE (Xiao et al., 2023) adopt a multi-stage training paradigm. They first pre-train on large-scale weakly-supervised text pairs using contrastive loss and then fine-tune on small-scale but high-quality datasets. In this paper, we demonstrate that it is possible to obtain state-of-the-art text embeddings with single-stage training.
Synthetic Data Synthetic data generation is a widely studied topic in information retrieval research, with various methods proposed to enhance retrieval systems with artificially created data. For instance, Doc2query (Nogueira et al., 2019), InPars (Bonifacio et al., 2022), and Promptagator (Dai et al., 2022) generate synthetic queries for unlabeled documents, which are then leveraged for document expansion or model training. GPL (Wang et al., 2022a) employs a cross-encoder to produce pseudo-labels for query-document pairs. Similarly, Query2doc (Wang et al., 2023) generates pseudo-documents for query expansion by few-shot prompting LLMs. Unlike these methods, our approach does not rely on any unlabeled documents or queries and thus can generate more diverse synthetic data.
Another related line of work focuses on knowledge distillation from black-box LLMs by training on synthetic data generated from them. DINO (Schick and Schütze, 2021) generates synthetic text pairs for semantic textual similarity. Unnatural Instructions (Honovich et al., 2022) is a synthetic instruction following dataset by prompting existing LLMs. Orca (Mukherjee et al., 2023) and Phi (Gunasekar et al., 2023) propose to train better small language models by using high-quality synthetic data from GPT-3.5/4 (OpenAI, 2023).
Large Language Models With the popularization of ChatGPT, large language models (LLMs) have demonstrated remarkable capabilities in instruction following and few-shot in-context learning (Brown et al., 2020). However, the most advanced LLMs such as GPT-4 (OpenAI, 2023) are proprietary and have little technical details disclosed. To bridge the gap between proprietary and open-source LLMs, several notable efforts have been made, such as LLaMA-2 (Touvron et al., 2023) and Mistral (Jiang et al., 2023) models. A major limitation of LLMs is that they lack awareness of recent events and private knowledge. This issue can be partly mitigated by augmenting LLMs with information retrieved from external sources, a technique known as retrieval-augmented generation (RAG). On the other hand, LLMs can also serve as foundation models to enhance text embeddings. RepLLaMA (Ma et al., 2023) proposes to fine-tune LLaMA-2 with bi-encoder architecture for ad-hoc retrieval. SGPT (Muennighoff, 2022), GTR (Ni et al., 2022b), and Udever (Zhang et al., 2023a) demonstrate the scaling law of text embeddings empirically, but their performance still falls behind small bidirectional encoders such as E5 (Wang et al., 2022b) and BGE (Xiao et al., 2023). In this paper, we present a novel approach to train state-of-the-art text embeddings by exploiting the latest advances of LLMs and synthetic data.
3 Method
3.1 Synthetic Data Generation
Utilizing synthetic data generated by advanced LLMs such as GPT-4 presents a compelling opportunity, especially in terms of enhancing diversity across a multitude of tasks and languages. Such diversity is essential for developing robust text embeddings that can perform well across different tasks, be it semantic retrieval, textual similarity, or clustering.
To generate diverse synthetic data, we propose a simple taxonomy that categorizes embedding tasks into several groups, and then apply different prompt templates to each group.
Asymmetric Tasks This category comprises tasks where the query and document are semantically related but are not paraphrases of each other. Depending on the length of the query and document, we further divide asymmetric tasks into four subgroups: short-long match, long-short match, short-short match, and long-long match. For instance, short-long match tasks involve a short query and a long document, which is a typical scenario in commercial search engines. For each subgroup, we design a two-step prompt template that first prompts LLMs brainstorm a list of tasks, and then generates a concrete example conditioned on the task definition. In Figure 1, we show an example prompt for the short-long match subgroup. The outputs from GPT-4 are mostly coherent and of high quality. In our preliminary experiments, we also attempted to generate the task definition and query-document pairs using a single prompt, but the data diversity was not as satisfactory as the proposed two-step approach.
Symmetric Tasks Symmetric tasks involve queries and documents that have similar semantic meanings but different surface forms. We examine two application scenarios: monolingual semantic textual similarity (STS) and bitext retrieval. We design two distinct prompt templates for each scenario, tailored to their specific objectives. Since the task definition is straightforward, we omit the brainstorming step for symmetric tasks.
To further boost the diversity of the prompts and thus the synthetic data, we incorporate several placeholders in each prompt template, whose values are randomly sampled at runtime. For example, in Figure 1, the value of “{query_length}” is sampled from the set “{less than 5 words, 5-10 words, at least 10 words}”.
To generate multilingual data, we sample the value of “{language}” from the language list of XLM-R (Conneau et al., 2020), giving more weight to high-resource languages. Any generated data that does not conform to the predefined JSON format are discarded during the parsing process. We also remove duplicates based on exact string matching.
3.2 Training
Given a relevant query-document pair (), we first apply the following instruction template to the original query to generate a new one :
where “{task_definition}” is a placeholder for a one-sentence description of the embedding task. For generated synthetic data, we use the outputs from the brainstorming step. For other datasets, such as MS-MARCO, we manually craft the task definitions and apply them to all the queries in the dataset. We do not modify the document side with any instruction prefix. In this way, the document index can be prebuilt, and we can customize the task to perform by changing only the query side.
Given a pretrained LLM, we append an [EOS] token to the end of the query and document, and then feed them into the LLM to obtain the query and document embeddings () by taking the last layer [EOS] vector. To train the embedding model, we adopt the standard InfoNCE loss over the in-batch negatives and hard negatives:
where denotes the set of all negatives, and is a function that computes the matching score between query and document . In this paper, we adopt the temperature-scaled cosine similarity function as follows:
is a temperature hyper-parameter, which is fixed to in our experiments.
4 Experiments
4.1 Statistics of the Synthetic Data
Figure 2 presents the statistics of our generated synthetic data. We manage to generate k examples with k unique instructions using Azure OpenAI Service 111https://oai.azure.com/, among which are generated by GPT-35-Turbo and others are generated by GPT-4. The total token consumption is about M. The predominant language is English, with coverage extending to a total of languages. For the bottom low-resource languages, there are about k examples per language on average.
In terms of data quality, we find that a portion of GPT-35-Turbo outputs do not strictly follow the guidelines specified in the prompt templates. Nevertheless, the overall quality remains acceptable, and preliminary experiments have demonstrated the benefits of incorporating this data subset.
4.2 Model Fine-tuning and Evaluation
The pretrained Mistral-7b (Jiang et al., 2023) checkpoint is fine-tuned for epoch using the loss in Equation 2. We follow the training recipe from RankLLaMA (Ma et al., 2023) and utilize LoRA (Hu et al., 2022) with rank . To further reduce GPU memory requirement, techniques including gradient checkpointing, mixed precision training, and DeepSpeed ZeRO-3 are applied.
For the training data, we utilize both the generated synthetic data and a collection of public datasets, yielding approximately M examples after sampling. More details are available in Appendix A. To provide a fair comparison with some previous work, we also report results when the only labeled supervision is the MS-MARCO passage ranking (Campos et al., 2016) dataset.
We evaluate the trained model on the MTEB benchmark (Muennighoff et al., 2023). Note that the retrieval category in MTEB corresponds to the publicly available datasets in the BEIR benchmark (Thakur et al., 2021). Evaluation of one model takes about days on V100 GPUs due to the need to encode a large number of documents. Although our model can accommodate sequence length beyond , we only evaluate on the first tokens for efficiency. Official metrics are reported for each category. For more details about the evaluation protocol, please refer to the original papers (Muennighoff et al., 2023; Thakur et al., 2021).
4.3 Main Results
| # of datasets | Class. | Clust. | PairClass. | Rerank | Retr. | STS | Summ. | Avg |
| 12 | 11 | 3 | 4 | 15 | 10 | 1 | 56 | |
| Unsupervised Models | ||||||||
| Glove (Pennington et al., 2014) | 57.3 | 27.7 | 70.9 | 43.3 | 21.6 | 61.9 | 28.9 | 42.0 |
| SimCSE (Gao et al., 2021) | 62.5 | 29.0 | 70.3 | 46.5 | 20.3 | 74.3 | 31.2 | 45.5 |
| Supervised Models | ||||||||
| SimCSE (Gao et al., 2021) | 67.3 | 33.4 | 73.7 | 47.5 | 21.8 | 79.1 | 23.3 | 48.7 |
| Contriever (Izacard et al., 2021) | 66.7 | 41.1 | 82.5 | 53.1 | 41.9 | 76.5 | 30.4 | 56.0 |
| GTR (Ni et al., 2022b) | 67.4 | 42.4 | 86.1 | 56.7 | 48.5 | 78.4 | 30.6 | 59.0 |
| Sentence-T5 (Ni et al., 2022a) | 73.4 | 43.7 | 85.1 | 56.4 | 42.2 | 82.6 | 30.1 | 59.5 |
| E5 (Wang et al., 2022b) | 75.2 | 44.5 | 86.0 | 56.6 | 50.6 | 82.1 | 30.2 | 62.3 |
| GTE (Li et al., 2023) | 73.3 | 46.8 | 85.0 | 59.1 | 52.2 | 83.4 | 31.7 | 63.1 |
| BGE (Xiao et al., 2023) | 76.0 | 46.1 | 87.1 | 60.0 | 54.3 | 83.1 | 31.6 | 64.2 |
| Ours | ||||||||
| E5 + full data | 78.5 | 50.3 | 88.3 | 60.2 | 56.9 | 84.6 | 31.4 | 66.6 |
| w/ synthetic data only | 78.2 | 50.5 | 86.0 | 59.0 | 46.9 | 81.2 | 31.9 | 63.1 |
| w/ synthetic + msmarco | 78.3 | 49.9 | 87.1 | 59.5 | 52.2 | 81.2 | 32.7 | 64.5 |
In Table 1, our model “E5 + full data” attains the highest average score on the MTEB benchmark, outperforming the previous state-of-the-art model by points. In the “w/ synthetic data only” setting, no labeled data is used for training, and yet the performance remains quite competitive. We posit that generative language modeling and text embeddings are the two sides of the same coin, with both tasks requiring the model to have a deep understanding of the natural language. Given an embedding task definition, a truly robust LLM should be able to generate training data on its own and then be transformed into an embedding model through light-weight fine-tuning. Our experiments shed light on the potential of this direction, and more research is needed to fully explore it.
| Model | BEIR Retrieval (15 datasets) | MTEB Average (56 datasets) |
| OpenAI Ada-002 | 49.3 | 61.0 |
| Cohere-embed-english-v3.0 | 55.0 | 64.5 |
| voyage-lite-01-instruct | 55.6 | 64.5 |
| UAE-Large-V1 (Li and Li, 2023) | 54.7 | 64.6 |
| E5 + full data | 56.9 | 66.6 |
In Table 2, we also present a comparison with several commercial text embedding models. However, due to the lack of transparency and documentation about these models, a fair comparison is not feasible. We focus especially on the retrieval performance on the BEIR benchmark, since RAG is an emerging technique to enhance LLM with external knowledge and proprietary data. As Table 2 shows, our model outperforms the current commercial models by a significant margin.
4.4 Multilingual Retrieval
| High-resource Languages | Low-resource Languages | |||||||
| en | fr | es | ru | te | hi | bn | sw | |
| BM25 (Zhang et al., 2023b) | 35.1 | 18.3 | 31.9 | 33.4 | 49.4 | 45.8 | 50.8 | 38.3 |
| mDPR (Zhang et al., 2023b) | 39.4 | 43.5 | 47.8 | 40.7 | 35.6 | 38.3 | 44.3 | 29.9 |
| mE5 (Wang et al., 2022b) | 51.2 | 49.7 | 51.5 | 61.5 | 75.2 | 58.4 | 70.2 | 71.1 |
| mE5 (Wang et al., 2022b) | 52.9 | 54.5 | 52.9 | 67.4 | 84.6 | 62.0 | 75.9 | 74.9 |
| E5 + full data | 57.3 | 55.2 | 52.2 | 67.7 | 73.9 | 52.1 | 70.3 | 68.4 |
To assess the multilingual capabilities of our model, we conduct an evaluation on the MIRACL dataset (Zhang et al., 2023b), which comprises human-annotated queries and relevance judgments across languages. As shown in Table 3, our model surpasses mE5 on high-resource languages, notably on English. Nevertheless, for low-resource languages, our model remains suboptimal compared to mE5. We attribute this to the fact that Mistral-7B is predominantly pre-trained on English data, and we anticipate that future multilingual LLMs will leverage our method to bridge this gap.
5 Analysis
5.1 Is Contrastive Pre-training Necessary?
Weakly-supervised contrastive pre-training is one of the key factors behind the success of existing text embedding models. For instance, Contriever (Izacard et al., 2021) treats random cropped spans as positive pairs for pre-training, while E5 (Wang et al., 2022b) and BGE (Xiao et al., 2023) collect and filter text pairs from various sources.
This section re-evaluates the necessity of contrastive pre-training for LLMs, particularly those that have been pre-trained on trillions of tokens. Figure 3 shows that contrastive pre-training benefits XLM-R, enhancing its retrieval performance by points when fine-tuned on the same data, which aligns with prior findings. However, for Mistral-7B based models, contrastive pre-training has negligible impact on the model quality. This implies that extensive auto-regressive pre-training enables LLMs to acquire good text representations, and only minimal fine-tuning is required to transform them into effective embedding models.
5.2 Extending to Long Text Embeddings
Existing evaluation datasets for text embedding models are typically short, to evaluate the long-context capability of our model, we introduce a novel synthetic task called personalized passkey retrieval, which is illustrated in Figure 4. This task requires encoding the passkey information in a long context into the embeddings. We compare the performance of different variants by changing the sliding window size and the RoPE rotation base (Su et al., 2024) in Figure 5. The results show that the default configuration with k sliding window attains accuracy within k tokens, but the accuracy deteriorates quickly as the context length grows. Naively extending the sliding window size to k results in worse performance. By changing the RoPE rotation base to , the model can achieve over accuracy within k tokens. However, this entails a minor trade-off in performance for shorter contexts. A potential avenue for future research is to efficiently adapt the model to longer contexts through lightweight post-training (Zhu et al., 2023).
5.3 Analysis of Training Hyperparameters
| Datasets | Class. | Clust. | PairClass. | Rerank | Retr. | STS | Summ. | Avg |
| E5 | 78.3 | 49.9 | 87.1 | 59.5 | 52.2 | 81.2 | 32.7 | 64.5 |
| w/ LLaMA-2 7b init. | 76.2 | 48.1 | 85.1 | 58.9 | 49.6 | 81.2 | 30.8 | 62.9 |
| w/ msmarco data only | 71.6 | 47.1 | 86.1 | 58.8 | 54.4 | 79.5 | 31.7 | 62.7 |
| pooling type | ||||||||
| w/ mean pool | 77.0 | 48.9 | 86.1 | 59.2 | 52.4 | 81.4 | 30.8 | 64.1 |
| w/ weighted mean | 77.0 | 49.0 | 86.1 | 59.2 | 52.0 | 81.4 | 30.2 | 64.0 |
| LoRA rank | ||||||||
| w/ r= | 78.4 | 50.3 | 87.1 | 59.3 | 53.0 | 81.0 | 31.7 | 64.8 |
| w/ r= | 78.4 | 50.3 | 87.4 | 59.5 | 52.2 | 81.2 | 30.6 | 64.6 |
| instruction type | ||||||||
| w/o instruction | 72.3 | 47.1 | 82.6 | 56.3 | 48.2 | 76.7 | 30.7 | 60.3 |
| w/ task type prefix | 71.1 | 46.5 | 79.7 | 54.0 | 52.7 | 73.8 | 30.0 | 60.3 |
Table 4 presents the results under different configurations. We notice that the Mistral-7B initialization holds an advantage over LLaMA-2 7B, in line with the findings from Mistral-7B technical report (Jiang et al., 2023). The choice of pooling types and LoRA ranks does not affect the overall performance substantially, hence we adhere to the default setting despite the marginal superiority of LoRA rank . On the other hand, the way of adding instructions has a considerable impact on the performance. We conjecture that natural language instructions better inform the model regarding the embedding task at hand, and thus enable the model to generate more discriminative embeddings. Our framework also provides a way to customize the behavior of text embeddings through instructions without the need to fine-tune the model or re-built document index.
6 Conclusion
This paper shows that the quality of text embeddings can be substantially enhanced by exploiting LLMs. We prompt proprietary LLMs such as GPT-4 to generate diverse synthetic data with instructions in many languages. Combined with the strong language understanding capability of the Mistral model, we establish new state-of-the-art results for nearly all task categories on the competitive MTEB benchmark. The training process is much more streamlined and efficient than existing multi-stage approaches, thereby obviating the need for intermediate pre-training.
For future work, we aim to further improve the multilingual performance of our model and explore the possibility of using open-source LLMs to generate synthetic data. We also intend to investigate ways to improve the inference efficiency and lower the storage cost for LLM based text embeddings.
Appendix A Implementation Details
Baseline Models For results with mE5 and mE5, we use the public checkpoints available at https://huggingface.co/intfloat/multilingual-e5-base and https://huggingface.co/intfloat/multilingual-e5-large respectively. For experiments in Table 4, we follow the SGPT [Muennighoff, 2022] paper for the implementation of weighted mean pooling. For the “w/ task type prefix” setting, we prepend “classify: ” for the long-short matching subgroup, and “query: ” for other asymmetric tasks. No prefix is added for symmetric tasks.
Training Data For the “E5 + full data” setting, our training data comprises generated synthetic data, ELI5 [Fan et al., 2019](sample ratio ), HotpotQA [Yang et al., 2018], FEVER [Thorne et al., 2018], MIRACL [Zhang et al., 2023b], MSMARCO passage ranking (sample ratio ) and document ranking (sample ratio ) [Campos et al., 2016], NQ [Karpukhin et al., 2020], NLI [Gao et al., 2021], SQuAD [Karpukhin et al., 2020], TriviaQA [Karpukhin et al., 2020], Quora Duplicate Questions [DataCanary et al., 2017](sample ratio ), MrTyDi [Zhang et al., 2021], DuReader [Qiu et al., 2022], and T2Ranking [Xie et al., 2023](sample ratio ) datasets. We only include the training set of each dataset. For the datasets without hard negatives, we use mE5 to mine top hard negatives. After sampling, we obtain approximately million examples. The entire training process takes fewer than k steps to complete.
Hyperparameters for Fine-tuning When fine-tuning Mistral-7b, the batch size is set to and the learning rate is with step warmup and linear decay. The weight decay is . We add hard negative for each query-document pair. The fine-tuning process takes roughly hours on V100 GPUs with a maximum sequence length . We add LoRA adapters to all linear layers, resulting in a total of M trainable parameters. Our implementation is based on the HuggingFace PEFT library at https://github.com/huggingface/peft.
The model and dataset release information is available at https://github.com/microsoft/unilm/tree/master/e5.
Appendix B Test Set Contamination Analysis
To assess the test set contamination on all the datasets in the MTEB benchmark, we perform a string match based analysis between the test set and our training set, disregarding differences in character case and spacing. We categorize the train-test overlaps into three types:
Low entropy texts. These are texts such as “i need a coffee” and “what does that mean”, which are not considered as contamination because they are common expressions that can occur in various contexts.
Question overlap. We identify test set questions in the DBPedia dataset that also appear in the TriviaQA training set. Given that they constitute a minor portion of the test set, their impact on the overall performance is insignificant.
Retrieval corpus overlap. Several retrieval datasets share the same retrieval corpus. For instance, the DBPedia, NQ, and TriviaQA datasets all use Wikipedia passages, even though their query sets are different. This is a standard evaluation practice in the field of information retrieval, and we do not regard it as contamination.
In summary, we did not detect substantial contamination risks that could alter the main findings of this paper.
Another aspect to consider is the possibility of test set contamination in the training data of Mistral-7B and GPT-4. However, since the training data of these models is not publicly accessible, it is challenging to estimate the degree of such contamination. Given their widespread use in the research community, we believe it is still a valid comparison if other works also employ these models.
| nDCG@10 | Recall@100 | |||||||||
| BM25 | mDPR | mE5 | mE5 | E5 full | BM25 | mDPR | mE5 | mE5 | E5 full | |
| ar | 48.1 | 49.9 | 71.6 | 76.0 | 73.3 | 88.9 | 84.1 | 95.9 | 97.3 | 96.0 |
| bn | 50.8 | 44.3 | 70.2 | 75.9 | 70.3 | 90.9 | 81.9 | 96.6 | 98.2 | 96.0 |
| en | 35.1 | 39.4 | 51.2 | 52.9 | 57.3 | 81.9 | 76.8 | 86.4 | 87.6 | 90.2 |
| es | 31.9 | 47.8 | 51.5 | 52.9 | 52.2 | 70.2 | 86.4 | 88.6 | 89.1 | 87.5 |
| fa | 33.3 | 48.0 | 57.4 | 59.0 | 52.1 | 73.1 | 89.8 | 91.2 | 92.9 | 88.0 |
| fi | 55.1 | 47.2 | 74.4 | 77.8 | 74.7 | 89.1 | 78.8 | 96.9 | 98.1 | 96.7 |
| fr | 18.3 | 43.5 | 49.7 | 54.5 | 55.2 | 65.3 | 91.5 | 90.0 | 90.6 | 92.8 |
| hi | 45.8 | 38.3 | 58.4 | 62.0 | 52.1 | 86.8 | 77.6 | 92.6 | 93.9 | 89.9 |
| id | 44.9 | 27.2 | 51.1 | 52.9 | 52.7 | 90.4 | 57.3 | 87.4 | 87.9 | 88.4 |
| ja | 36.9 | 43.9 | 64.7 | 70.6 | 66.8 | 80.5 | 82.5 | 96.0 | 97.1 | 95.1 |
| ko | 41.9 | 41.9 | 62.2 | 66.5 | 61.8 | 78.3 | 73.7 | 91.6 | 93.4 | 89.4 |
| ru | 33.4 | 40.7 | 61.5 | 67.4 | 67.7 | 66.1 | 79.7 | 92.7 | 95.5 | 95.0 |
| sw | 38.3 | 29.9 | 71.1 | 74.9 | 68.4 | 70.1 | 61.6 | 95.6 | 96.7 | 95.5 |
| te | 49.4 | 35.6 | 75.2 | 84.6 | 73.9 | 83.1 | 76.2 | 98.0 | 99.2 | 95.1 |
| th | 48.4 | 35.8 | 75.2 | 80.2 | 74.0 | 88.7 | 67.8 | 98.0 | 98.9 | 96.5 |
| zh | 18.0 | 51.2 | 51.5 | 56.0 | 54.0 | 56.0 | 94.4 | 92.1 | 93.3 | 90.1 |
| Avg | 39.3 | 41.5 | 62.3 | 66.5 | 62.9 | 78.7 | 78.8 | 93.1 | 94.3 | 92.6 |
| Datasets | Class. | Clust. | PairClass. | Rerank | Retr. | STS | Summ. | Avg |
| XLM-R + full data | 72.9 | 38.7 | 84.5 | 53.8 | 42.0 | 82.3 | 29.7 | 58.0 |
| w/ cont. pre-train | 77.2 | 47.3 | 85.5 | 58.6 | 50.2 | 84.4 | 30.7 | 63.7 |
| E5 + full data | 78.5 | 50.3 | 88.3 | 60.2 | 56.9 | 84.6 | 31.4 | 66.6 |
| w/ cont. pre-train | 78.7 | 50.1 | 87.7 | 60.9 | 56.9 | 84.9 | 30.2 | 66.7 |
Appendix C Prompts for Synthetic Data Generation
For asymmetric tasks, we list the four prompt templates in Table 7, 8, 9, and 10. For symmetric tasks, the prompts templates are available in Table 11 and 12. To generate multilingual data, we sample the value of “{language}” from the language list of XLM-R [Conneau et al., 2020] with higher probability for high-resource languages. When prompting GPT-4/3.5, we set the temperature to and the top- to , which is higher than the default setting to encourage more diversity.
Appendix D Instructions for Training and Evaluation
| Dataset | Instruction | ||||
| ELI5 |
| HotpotQA | Given a multi-hop question, retrieve documents that can help answer the question | ||||
| FEVER | Given a claim, retrieve documents that support or refute the claim | ||||
| NLI |
| MS-MARCO |
| Quora Duplicates |
| DuReader / T2Ranking | Given a Chinese search query, retrieve web passages that answer the question |
| Task Name | Instruction | |
| AmazonCounterfactualClassif. |
| AmazonPolarityClassification | Classify Amazon reviews into positive or negative sentiment | |
| AmazonReviewsClassification | Classify the given Amazon review into its appropriate rating category | |
| Banking77Classification | Given a online banking query, find the corresponding intents | |
| EmotionClassification |
| ImdbClassification | Classify the sentiment expressed in the given movie review text from the IMDB dataset | |
| MassiveIntentClassification | Given a user utterance as query, find the user intents | |
| MassiveScenarioClassification | Given a user utterance as query, find the user scenarios | |
| MTOPDomainClassification | Classify the intent domain of the given utterance in task-oriented conversation | |
| MTOPIntentClassification | Classify the intent of the given utterance in task-oriented conversation | |
| ToxicConversationsClassif. | Classify the given comments as either toxic or not toxic | |
| TweetSentimentClassification | Classify the sentiment of a given tweet as either positive, negative, or neutral | |
| ArxivClusteringP2P |
| ArxivClusteringS2S | Identify the main and secondary category of Arxiv papers based on the titles | |
| BiorxivClusteringP2P | Identify the main category of Biorxiv papers based on the titles and abstracts | |
| BiorxivClusteringS2S | Identify the main category of Biorxiv papers based on the titles | |
| MedrxivClusteringP2P | Identify the main category of Medrxiv papers based on the titles and abstracts | |
| MedrxivClusteringS2S | Identify the main category of Medrxiv papers based on the titles | |
| RedditClustering | Identify the topic or theme of Reddit posts based on the titles | |
| RedditClusteringP2P | Identify the topic or theme of Reddit posts based on the titles and posts | |
| StackExchangeClustering | Identify the topic or theme of StackExchange posts based on the titles | |
| StackExchangeClusteringP2P | Identify the topic or theme of StackExchange posts based on the given paragraphs | |
| TwentyNewsgroupsClustering | Identify the topic or theme of the given news articles | |
| SprintDuplicateQuestions | Retrieve duplicate questions from Sprint forum | |
| TwitterSemEval2015 | Retrieve tweets that are semantically similar to the given tweet | |
| TwitterURLCorpus | Retrieve tweets that are semantically similar to the given tweet | |
| AskUbuntuDupQuestions | Retrieve duplicate questions from AskUbuntu forum | |
| MindSmallReranking | Retrieve relevant news articles based on user browsing history | |
| SciDocsRR | Given a title of a scientific paper, retrieve the titles of other relevant papers | |
| StackOverflowDupQuestions | Retrieve duplicate questions from StackOverflow forum | |
| ArguAna | Given a claim, find documents that refute the claim | |
| ClimateFEVER | Given a claim about climate change, retrieve documents that support or refute the claim | |
| CQADupstackRetrieval |
| DBPedia | Given a query, retrieve relevant entity descriptions from DBPedia | |
| FEVER | Given a claim, retrieve documents that support or refute the claim | |
| FiQA2018 | Given a financial question, retrieve user replies that best answer the question | |
| HotpotQA | Given a multi-hop question, retrieve documents that can help answer the question | |
| MSMARCO | Given a web search query, retrieve relevant passages that answer the query | |
| NFCorpus | Given a question, retrieve relevant documents that best answer the question | |
| NQ | Given a question, retrieve Wikipedia passages that answer the question | |
| QuoraRetrieval |
| SCIDOCS | Given a scientific paper title, retrieve paper abstracts that are cited by the given paper | |
| SciFact | Given a scientific claim, retrieve documents that support or refute the claim | |
| Touche2020 | Given a question, retrieve detailed and persuasive arguments that answer the question | |
| TRECCOVID | Given a query on COVID-19, retrieve documents that answer the query | |
| STS* | Retrieve semantically similar text. | |
| BUCC/Tatoeba | Retrieve parallel sentences. | |
| SummEval | Given a news summary, retrieve other semantically similar summaries |
We manually write instructions for training datasets, as listed in Table 13. For evaluation datasets, the instructions are listed in Table 14.
| Dataset | w/ synthetic only | w/ synthetic + msmarco | w/o synthetic data | full data |
| BIOSSES | 84.2 | 81.0 | 85.4 | 85.5 |
| SICK-R | 78.6 | 78.5 | 81.7 | 82.6 |
| STS12 | 75.8 | 74.7 | 77.9 | 79.7 |
| STS13 | 84.3 | 85.3 | 88.0 | 88.4 |
| STS14 | 80.9 | 81.2 | 83.7 | 84.5 |
| STS15 | 86.2 | 86.8 | 89.5 | 90.4 |
| STS16 | 85.0 | 85.3 | 86.5 | 87.7 |
| STS17 | 87.3 | 87.7 | 91.0 | 91.8 |
| STS22 | 66.0 | 67.1 | 66.2 | 67.0 |
| STSBenchmark | 83.5 | 84.0 | 87.8 | 88.6 |
| SummEval | 31.9 | 32.7 | 31.9 | 31.4 |
| SprintDuplicateQuestions | 93.5 | 95.8 | 96.0 | 95.7 |
| TwitterSemEval2015 | 78.0 | 78.5 | 81.7 | 81.6 |
| TwitterURLCorpus | 86.5 | 86.9 | 87.7 | 87.8 |
| AmazonCounterfactualClass. | 79.6 | 79.9 | 77.2 | 78.7 |
| AmazonPolarityClassification | 95.8 | 95.9 | 93.9 | 95.9 |
| AmazonReviewsClassification | 56.9 | 55.5 | 48.2 | 55.8 |
| Banking77Classification | 86.2 | 87.0 | 88.8 | 88.2 |
| EmotionClassification | 49.2 | 47.6 | 51.0 | 49.8 |
| ImdbClassification | 94.8 | 94.9 | 89.0 | 94.8 |
| MassiveIntentClassification | 79.8 | 79.9 | 79.6 | 80.6 |
| MassiveScenarioClassification | 81.7 | 82.4 | 82.3 | 82.4 |
| MTOPDomainClassification | 95.6 | 95.9 | 95.7 | 96.1 |
| MTOPIntentClassification | 84.9 | 85.9 | 83.4 | 86.1 |
| ToxicConversationsClassification | 70.2 | 70.8 | 70.9 | 69.6 |
| TweetSentimentExtractionClass. | 63.5 | 63.4 | 61.6 | 63.7 |
| AskUbuntuDupQuestions | 64.3 | 65.3 | 67.4 | 67.0 |
| MindSmallReranking | 33.1 | 32.8 | 32.5 | 32.6 |
| SciDocsRR | 86.0 | 86.0 | 85.7 | 86.3 |
| StackOverflowDupQuestions | 52.5 | 53.7 | 55.9 | 54.9 |
| ArxivClusteringP2P | 51.4 | 51.2 | 47.8 | 50.5 |
| ArxivClusteringS2S | 46.5 | 44.9 | 44.6 | 45.5 |
| BiorxivClusteringP2P | 44.5 | 43.3 | 36.9 | 43.5 |
| BiorxivClusteringS2S | 40.9 | 40.1 | 37.0 | 40.2 |
| MedrxivClusteringP2P | 40.5 | 39.9 | 32.6 | 38.2 |
| MedrxivClusteringS2S | 38.0 | 37.9 | 32.8 | 37.5 |
| RedditClustering | 56.3 | 55.9 | 63.1 | 57.7 |
| RedditClusteringP2P | 66.3 | 64.8 | 66.4 | 66.5 |
| StackExchangeClustering | 72.9 | 72.7 | 74.5 | 73.1 |
| StackExchangeClusteringP2P | 46.1 | 45.6 | 34.3 | 45.9 |
| TwentyNewsgroupsClustering | 52.2 | 52.5 | 55.6 | 54.3 |
| ArguAna | 52.2 | 42.7 | 62.5 | 61.9 |
| ClimateFEVER | 21.1 | 28.8 | 25.2 | 38.4 |
| CQADupstackAndroidRetrieval | 40.8 | 36.0 | 44.5 | 43.0 |
| DBPedia | 42.0 | 43.7 | 47.7 | 48.9 |
| FEVER | 72.5 | 83.5 | 73.1 | 87.8 |
| FiQA2018 | 38.1 | 48.4 | 54.5 | 56.6 |
| HotpotQA | 48.1 | 64.0 | 75.6 | 75.7 |
| MSMARCO | 25.7 | 45.0 | 42.9 | 43.1 |
| NFCorpus | 35.5 | 40.0 | 35.3 | 38.6 |
| NQ | 53.3 | 63.5 | 57.3 | 63.5 |
| QuoraRetrieval | 75.0 | 79.5 | 89.5 | 89.6 |
| SCIDOCS | 20.6 | 15.8 | 19.0 | 16.3 |
| SciFact | 71.5 | 71.9 | 74.7 | 76.4 |
| Touche2020 | 25.4 | 32.5 | 19.1 | 26.4 |
| TRECCOVID | 82.3 | 87.3 | 70.8 | 87.2 |
| Average | 63.1 | 64.5 | 64.6 | 66.6 |